(1) A 31.4 mL sample of a 0.531 M aqueous nitrous acid solution is titrated with a 0.305 M aqueous barium hydroxide solution. What is the pH after 10.4 mL of base have been added?

(2) When a 28.7 mL sample of a 0.424 M aqueous hydrofluoric acid solution is titrated with a 0.352 M aqueous sodium hydroxide solution, what is the pH after 51.9 mL of sodium hydroxide have been added?
Answer #1

. Kao HNO2 = 4.5 X10-4 pko = -log kq = -109 (4.5 X10-4 ) = 3.35 milimole immol) - concentration (M) X volumeamL). mmol of HNO[NO₂-] = 0.15178 M equation => [HNO, ] = 0.24712 M using Hendeuon Haralbalch ÞHe þke + 109 [NO2-] (HNO3) pH = 3.35 + log (6.9kindly post different questions separately. please rate